myininaya (myininaya):

Each employee of a certain company is in either Department X or Department Y , and there are more than twice as many employees in Department X as in Department Y. The average salary is 25,000 dollars for the employees in Department X and is 35,000 for the employees in Department Y. Which of the following amounts could be average salary for all of the employees in the company? Include all such amounts. A. 26,000 B. 28,000 C. 29,000 D.30,000 E. 31,000 F. 32,000 G. 34,000

OpenStudy (anonymous):

myin?

myininaya (myininaya):

hey

OpenStudy (anonymous):

I guess this would be weighted average?

OpenStudy (anonymous):

how was calc II at dawn?

myininaya (myininaya):

lol it was okay

OpenStudy (phi):

Does this make sense: let y= # in dept Y. assume # in dept X is exactly 2y then total salary = 2y*25 + y*35 and avg salary would be 85y/3y = 28.33 so moving people from dept Y to dept X would lower this to 28. In the limit, if all people moved from Y to X, the avg would be 25, so 26 is also possible

hero (hero):

Hi Apple Pi

OpenStudy (anonymous):

i was thinking that you have average is \[\frac{25x+35y}{x+y}\] with the obvious meaning of x and y. also you know \[x>2y\] so smallest denominator could be is \[3y\] if x =2 y, and biggest numerator could be is also if x = 2y at \[95y\] so biggest whole fraction can be is \[\frac{95y}{3y}=31\tfrac{2}{3}\] so anything less

OpenStudy (anonymous):

smallest is if they all work in X right? in that case you get \[25\]

OpenStudy (anonymous):

but i have been making mistakes all day. so could be wrong

myininaya (myininaya):

phi good job thats what i did A and B are the answers since we have that or situation

OpenStudy (phi):

@sat if you sub in 2y for x, you get 50y+35y. It looks like you did it the other way round. BTW, how do you get your post to be formatted??

OpenStudy (anonymous):

told you i had made mistakes all day!

myininaya (myininaya):

lol

myininaya (myininaya):

phi are you asking how to type in latex?

myininaya (myininaya):

satellite they aren't mistakes they are type-o's remember?